# Repositioning, Ride-matching, and Abandonment in On-demand Ride-hailing Platforms: A Mean Field Game Approach

### v1

@@ Abstract.
The on-demand ride-hailing industry has experienced rapid growth, transforming transportation norms worldwide. Despite improvements in efficiency over traditional taxi services, significant challenges remain, including drivers’ strategic repositioning behavior, customer abandonment, and inefficiencies in dispatch algorithms. To address these issues, we introduce a comprehensive mean field game model that systematically analyzes the dynamics of ride-hailing platforms by incorporating driver repositioning across multiple regions, customer abandonment behavior, and platform dispatch algorithms. Using this framework, we identify all possible mean field equilibria as the Karush-Kuhn-Tucker (KKT) points of an associated optimization problem. Our analysis reveals the emergence of multiple equilibria, including the inefficient “Wild Goose Chase” one, characterized by drivers pursuing distant requests, leading to suboptimal system performance. To mitigate these inefficiencies, we propose a novel two-matching-radius nearest-neighbor dispatch algorithm that eliminates undesirable equilibria and ensures a unique mean field equilibrium for multi-region systems. The algorithm dynamically adjusts matching radii based on driver supply rates, optimizing pick-up times and waiting times for drivers while maximizing request completion rates. Numerical experiments and simulation results show that our proposed algorithm reduces customer abandonment, minimizes waiting times for both customers and drivers, and improves overall platform efficiency.
@@ 1. Introduction
The on-demand ride-hailing industry has undergone significant expansion in recent years. Uber’s span to over 700 cities worldwide and the pervasive presence of ride-hailing services in China signifies a remarkable shift in transportation norms for people. Although ride-hailing platforms improve efficiency over traditional taxi services, they still encounter challenges and opportunities for enhancement.
@@
Finally, the design of efficient dispatch algorithms is paramount to preventing the “Wild Goose Chase” (WGC) equilibrium (Castillo et al., 2023), where drivers chase distant requests inefficiently, leading to increased empty driving time and decreased overall system efficiency. Traditional matching algorithms may induce non-monotonic sojourn times in the driver supply rate, resulting in multiple equilibria and suboptimal outcomes. Innovating in this area requires the development of new algorithms that can dynamically adjust to real-time conditions and optimize the matching process to minimize inefficiencies and maximize service quality.
@@
To model the strategic behavior among a large number of agents (drivers), our study employs the framework of mean field games. This approach simplifies the complex interactions between numerous agents into a representative agent interacting with the mean field, thereby significantly reducing computational complexity. Moreover, when the number of agents is large, this approximation maintains a high degree of accuracy. Given the intricate operational and strategical details involving the platform, passengers, and drivers, constructing an exact model is exceedingly difficult, and precise analysis of such a complex model is nearly impossible. Therefore, we focus on a fundamental mean field game model that captures the core issues. Adopting an agent-based modeling perspective, we consider a closed queueing network comprising a large number of drivers, where each driver’s sojourn time follows a general distribution. The game model we propose is a continuous-time mean field game with finite state-action space where each driver’s decision model is a semi-Markov decision model (semi-MDP). Drawing inspiration from (Kelly, 1989), we apply mean field approximation to express the average sojourn times as functions of the system’s throughput, which in our model corresponds to the driver supply rate—that is, the number of available ride-hailing vehicles per unit time in a specified region. This allows us to transform the problem of finding mean field game equilibria into finding Karush-Kuhn-Tucker (KKT) points for optimization problems. This pivotal idea, initially applied to mean field games in (Courcoubetis and Dimakis, 2023), is further generalized and extended in our work.
@@
Employing the aforementioned mean field game model, this study examines the impact of customer abandonment behavior on platform dispatch algorithms. Specifically, we delve into the classical nearest-neighbor dispatch algorithm, which pairs the closest available vehicle with a customer in need. Our analysis reveals a significant WGC phenomenon. This phenomenon manifests as the existence of multiple equilibria within the system, where “bad equilibrium” results in longer pick-up times compared to “good equilibrium”. In the “bad equilibrium”, distant vehicles are matched with passengers, leading to substantial efficiency losses in the system, which theoretically could become infinitely large. Meanwhile, our analysis reveals that multiple equilibria emerge because the mean pick-up time is a non-monotone function of the driver supply rate. To mitigate the WGC issue, we investigate the effects of introducing matching radius—the maximum distance within which a match is allowed. Our findings indicate that the careful selection of the matching radius is crucial, as an improper choice could still result in multiple equilibria and the persistence of WGC scenarios. The presence of multiple equilibria complicates the task of selecting an optimal matching radius, as there is no explicit solution available. This selection process requires a nuanced consideration of various parameters, particularly the fluctuating number of vehicles in the system.
@@
In light of the difficulties associated with directly determining the optimal matching radius, our study innovatively proposes the Two-Matching-Radius Nearest-Neighbor Dispatch Algorithm, which can be viewed as a generalization of the traditional nearest-neighbor dispatch algorithm. This novel approach designs distinct matching radii for two scenarios: the arrival of new customers and the arrival of new available vehicles. Rather than seeking static optimal matching radii for a fixed system, we solve for optimal adjustment of matching radii, wherein for each given driver supply rate, we find the corresponding optimal matching radii. We derive an explicit relationship between the optimal matching radii and the driver supply rate. Furthermore, we demonstrate that the nearest-neighbor dispatch Algorithm utilizing the optimal adjustment of matching radii results in a sojourn time that is a monotone function of the driver supply rate. This allows us to prove the uniqueness of the mean field equilibrium, thereby theoretically eliminating the WGC equilibrium. By minimizing the sojourn time for each given driver supply rate through our optimal matching radius, our solution not only eradicates undesirable equilibria but also maximizes the system’s throughput, thereby optimizing overall system efficiency for ride-hailing platforms.
@@
  • We introduce a holistic mean field game model to systematically analyze the ride-hailing systems, incorporating key aspects such as driver’s repositioning across multiple regions, customer’s abandonment behavior, and platform’s dispatch algorithms. This unified model allows for the study of their joint effects on system performance.
  • In this general multi-region mean field game framework, we identify all possible mean field equilibria as the Karush-Kuhn-Tucker (KKT) points of an associated optimization problem. Our model reveals the emergence of multiple equilibria, including the inefficient “Wild Goose Chase” one – a well-known matching failure in ride-hailing systems extensively discussed in the literature (Castillo et al., 2023). In this WGC equilibrium, drivers are matched with distant requests and consequently the system efficiency is compromised.
  • We propose a simple yet innovative two-matching-radius nearest-neighbor dispatch algorithm. This algorithm eliminates undesirable equilibria, including the “Wild Goose Chase” one, and ensures a unique mean field equilibrium for the multi-region model.
@@ 1.1. Literature Review
Repositioning in Ride-hailing Platforms. (Braverman et al., 2019) studied the repositioning of empty cars in a spatial ride-sharing network where a platform controls the repositioning of all vehicles. (Afèche et al., 2023) studied strategical repositioning in ride-hailing networks where drivers are allowed to make repositioning decisions to maximize their earnings. (Benjaafar et al., 2021) and (Gao et al., 2024) considered a combination of decentralized and centralized repositioning to investigate the impact of autonomous vehicles on ride-hailing platforms. Our work is closely related to (Gao et al., 2024) in that we similarly consider a mean field game model and characterize the mean field equilibrium by an optimization problem. The matching procedure are simplified in these literature as they all explicitly or implicitly assume that rider requests are lost if not matched instantly with an available driver. In our model, a detailed matching procedure between riders and drivers is incorporated to the study of the strategical repositioning behavior of drivers.
@@
Ride-matching and Abandonment. Another stream of research close to our work is the study of the performance of matching in ride-hailing systems. (Besbes et al., 2022) considers nearest-neighbor dispatch and uses a spatial queueing model to derive the service rate as a function of pick time and En route time. (Wang et al., 2022) extends the spatial model in (Besbes et al., 2022) to include admission control for matching based on a pickup-time threshold. In addition, abandonment and cancellation on customer’s side are investigated in (Wang et al., 2022). Our model inherits the customer’s abandonment behavior model from (Wang et al., 2022) and similarly consider a two-sided queueing model with open customer-side queue and a closed driver-side queueing network. (Castillo et al., 2023) investigated the “Wild Goose Chase” caused by long pick-up times. (Castillo et al., 2023) suggested that setting a maximum matching radius could improve the system performance. (Yang et al., 2020) studied the impact of matching radius and matching time interval on the system performance. (Wang et al., 2022) and (Yang et al., 2020) both considered batch matching while (Wang et al., 2022) proposed to manipulate the matching radius and matching time interval indirectly through a pickup-time threshold. Different from (Wang et al., 2022) and (Yang et al., 2020), we consider instant matching using nearest-neighbor dispatch, i.e., an arriving customer is immediately matched with the closest idle driver and an arriving driver is immediately matched with the closest waiting customer. Thus, our work views matching radius as the key decision variable of the platform. Our research contribute to this stream of research by introducing a novel two-matching-radius nearest-neighbor dispatch algorithm. The dynamically optimal adjustment of the matching radii proposed in this work offers an solution to the WGC problem.
@@
Mean Field Game. Another area closely related to our work is mean-field games, which investigate the repeated interactions among a large number of strategic agents ((Huang et al., 2006), (Lasry and Lions, 2007)). Our mean-field game model is of finite type, featuring finite state and action spaces. Existing finite mean-field game models in the literature are typically based on either discrete-time Markov Decision Processes (MDPs) (Guo et al., 2023) or continuous-time MDPs with exponentially distributed sojourn times (Neumann, 2020). In our model, we model each driver’s decision-making by continuous-time semi-Markov decision processes (SMDPs) where sojourn times follow general distributions. In fact, our model incorporates a closed queueing network, where drivers can be viewed as “jobs” navigating through the network. Drawing inspiration from (Kelly, 1989), we express sojourn times as functions of driver supply rates. Building on this, we employ the methodology outlined in (Courcoubetis and Dimakis, 2023) to formulate an optimization problem for computing mean-field equilibria. However, the intricate nature of the closed spatial queueing network inherent in ride-hailing systems generally results in sojourn times that are not monotone functions of throughput, thereby violating the monontone assumptions put forth in (Kelly, 1989) and (Courcoubetis and Dimakis, 2023). To address this, we extend their methodologies to accommodate non-monotone scenarios. Moreover, by optimally designing the dispatch algorithm and matching radius, we can ensure that sojourn times become monotone, thereby validating the monotone assumptions of (Kelly, 1989) and (Courcoubetis and Dimakis, 2023).
@@ 1.2. Paper Organization
In Section 2, we describe each driver’s continuous-time semi-Markov Decision Process model and formulate a mean field game model. In Section 3, we investigate the ride-matching algorithm under abandonment behavior, derive the sojourn times as exact functions of driver supply rates. In Section 4, optimal matching radius design and a solution to eliminate the “Wild Goose Chase” equilibrium are provided . In Section 5, we conduct simulations to evaluate the performance our proposed dispatch algorithm. We conclude the paper in Section 6.
@@ 2.1. A Semi-Markov Decision Process
We consider a closed network comprising a total mass $m$ of drivers and $N$ regions, each region labeled by numbers $1,2,\cdots,N$. These regions can be viewed as distinct geographic areas within a city. The drivers are modeled by a continuous-time semi-Markov decision process (semi-MDP). The state space, denoted by $S=\{1,2,\cdots,N\}$, signifies the driver’s current location. The action space, denoted by $A=\{1,2,\cdots,N\}$ , indicates the driver’s intended destination. Hence, the semi-MDP we consider is a special one with $S=A$. For a given state-action pair $(i,k)\in S\times A=S^{2}$, $k=i$ implies that the driver chooses to provide services within region $i$, while $k\not=i$ indicates that the driver is repositioning to region $k$.
@@
Transition Ptobability. The transition probabilities are encapsulated in the tensor $P=(P^{k}_{ij})_{i,j\in S,k\in A}$, where the subscripts denote the state transition pairs and the superscript indicates the action. $P^{k}_{ij}$ denotes the probability of transitioning from state $i$ to state $j$ under action $k$. Let $q_{ik}$ denote the probability that a passenger in region $i$ has a destination in region $k$. Thus, the matrix $(q_{ik})_{(i,k)\in S\times A}$ is also referred to as the demand matrix. Using the demand matrix, we are able to define the transition matrix of the semi-MDP. The transition probability from state $i$ to state $j$ for any driver with the state-action pair $(i,k)$ is $P_{ij}^{k}=\delta_{jk}$ when $i\not=k$, where $\delta_{jk}=1$ if $j=k$ and 00 otherwise. For the state-action pair $(i,i)$, the transition probability from state $i$ to state $j$ is $P_{ij}^{k}=q_{ik}$. In this context, for a driver in state $i$ taking action $k$ where $k\not=i$, the driver opts to reposition to region $k$, leading to a state transition to region $k$ with probability one. When $k=i$, the driver decides to provide services within the current region, and the state transitions in accordance with the passenger demand matrix. In summary, the transition probabilities $(P^{k}_{ij})_{i,j\in S,k\in A}$ are given by,
$P_{ij}^{k}=\begin{cases}q_{ik}&\text{ if }i=j,\\ \delta_{jk}&\text{ if }i\not=j.\end{cases}$
@@
Sojourn Time. For state-action pair $(i,k)$, it takes a certain amount of time for a driver to complete action $k$ when they are in state/location $i$ (to serve a customer in region $i$ when $k=i$ or reposition to $k$ if $k\not=i$ ). This is also referred to as the sojourn time for the state-action pair $(i,k)$. Let $T_{ik}$ denote the mean time taken for a driver to finish action $k$ from state/location $i$. The sojourn time is a result of interactions across the platform, customers, and drivers, encompassing elements such as the driver’s wait time until a match, the duration of pick-ups, and travel times between locations. Factors influencing the determination of sojourn time include customer behavior, platform matching policies, and drivers’ strategies. Under appropriate assumptions and within specific customer behavior models and platform matching policies, we will show that the mean sojourn time can be calculated based on the distribution of driver masses over the state-action space and their respective mass rates.
@@
Steady State Analysis. We consider a non-atomic game model where a representative driver makes decisions against the mass behavior of all other drivers, characterized by the mass distribution over $S\times A$. This representative driver is negligible and exerts no influence on the system, including the mass distribution and corresponding mass rate. This approach is reasonable given the enormous number of drivers in the system, their anonymity, symmetry, and the sole differentiation based on their locations. We focus on stationary equilibria and study stationary systems, which is sensible since each driver perceives the system to be in a steady state, enabling observation of mass distributions and estimation of sojourn times effectively. Our approach involves a fluid-based analysis that operates with the averages of the sojourn times and associated metrics, a methodology commonly employed in related literature (Braverman et al., 2019; Besbes et al., 2022; Wang et al., 2022; Courcoubetis and Dimakis, 2023).
@@
Now, let us proceed to describe how to determine the sojourn times in a stationary system. The mean sojourn time for any $(i,k)\in S\times A$ is firstly dependent on the mean travel time $t_{ik}$ of a requested trip from region $i$ to region $k$ (similarly, $t_{ii}$ is the mean travel time for a requested trip within the region $i$). When $k\not=i$, the vehicle repositions and the mean sojourn time $T_{ik}$ equals the mean travel time $t_{ik}$. When $k=i$, the sojourn time consists of three components: waiting time to get matched to a customer, pick-up time and en route time, with means denoted by $w_{i}^{d}$, $\tau_{i}$ and $t_{i}$ respectively. The mean en route time $t_{i}$ can be computed from the demand matrix and mean travel time matrix $(t_{ij})_{i,j\in S}$ by $t_{i}=\sum_{j\in S}t_{ij}q_{ij}$. The mean waiting time of a driver $w_{i}^{d}$ and the mean pickup $\tau_{i}$ are influenced by both the customers’ arrival process and the platform’s matching policy. We will investigate these factors in the subsequent subsection. In fact, we can conceptualize that, based on the customer behavior model and platform matching policy, we are able to establish in each region relationships between the driver’s average waiting time, pick-up time, and mass rate. More specifically, the following assumption (that will be validated by the analysis) simply postulates that we can derive all other quantities as functions of the mass rate in each region. We use this assumption when we define the underlying mean-field game.
@@
For a given vector of customer arrival rates and platform matching policy, when the system is in a stationary state, the driver’s mean waiting time $w_{i}$ and mean pick-up time $\tau_{i}$ are positive and can be written as functions of the mass rate of available drivers $x_{ii}$, for every $i\in S$. Thus, the mean total sojourn time $T_{ii}$ can also be written as a function of $x_{ii}$, i.e.,
@@
where $t_{i}$ is a constant such that $t_{i}=\sum_{j\in S}t_{ij}q_{ij}$. Moreover, we further assume that as $x_{ii}\rightarrow b_{i}$, $w_{i}^{d}(x_{ii})\rightarrow\infty$.
It will be demonstrated that for the matching mechanism and customer abandonment behavior considered in this study, this assumption holds true. Explicit formulae will be provided to support this assumption. The assumption that as $x_{ii}\rightarrow b_{i}$, $w_{i}^{d}(x_{ii})\rightarrow\infty$ arises from the customer behavior model, where customers exhibit impatience and will abandon the service if not served promptly. As a result, an infinite number of idle drivers is necessary to meet all customer demands. This aspect will be elaborated on further in subsequent sections. In what follows, for the sake of uniformity, we write the sojourn time as a function of $x$, i.e., $T_{ik}(x)$ for any $(i,k)\in S\times A$, even though it is a function of $x_{ik}$ if $i=k$ and a constant if $i\not=k$. Given Assumption 1, the mass rate satisfies the following equation for a stationary system,

### v2

@@ Abstract.
The on-demand ride-hailing industry has experienced rapid growth, transforming transportation norms worldwide. Despite improvements in efficiency over traditional taxi services, significant challenges remain, including drivers’ strategic repositioning behavior, customer abandonment, and inefficiencies in dispatch algorithms. To address these issues, we introduce a comprehensive mean field game model that systematically analyzes the dynamics of ride-hailing platforms by incorporating driver repositioning across multiple regions, customer abandonment behavior, and platform dispatch algorithms. Using this framework, we identify all possible mean field equilibria as the Karush-Kuhn-Tucker (KKT) points of an associated optimization problem. Our analysis reveals the emergence of multiple equilibria, some of them characterized by drivers pursuing distant requests, i.e., exhibiting ‘Wild Goose Chase’ behavior, leading to suboptimal system performance. To mitigate these inefficiencies, we propose a novel two-matching-radius nearest-neighbor dispatch algorithm that eliminates undesirable equilibria and ensures a unique mean field equilibrium for multi-region systems. The algorithm dynamically adjusts matching radii based on driver supply rates, optimizing pick-up times and waiting times for drivers while maximizing request completion rates. Numerical experiments and simulation results show that our proposed algorithm reduces customer abandonment, minimizes waiting times for both customers and drivers, and improves overall platform efficiency.
@@ 1. Introduction
The on-demand ride-hailing industry has undergone significant expansion in recent years. Uber’s span to over 700 cities worldwide and the pervasive presence of ride-hailing services in China signify a remarkable shift in transportation norms for people. Although ride-hailing platforms improve efficiency over traditional taxi services, they still encounter challenges and opportunities for enhancement.
@@
Finally, the design of efficient dispatch algorithms is paramount to preventing Wild Goose Chase (WGC) equilibria (Castillo et al., 2023), where drivers chase distant requests inefficiently, leading to increased empty driving time and decreased overall system efficiency. Traditional matching algorithms may induce non-monotonic sojourn times in the driver supply rate, resulting in multiple equilibria and suboptimal outcomes. Innovating in this area requires the development of new algorithms that can dynamically adjust to real-time conditions and optimize the matching process to minimize inefficiencies and maximize service quality.
@@
To model the strategic behavior among a large number of agents (drivers), our study employs the framework of mean field games. This approach simplifies the complex interactions between numerous agents into a representative agent interacting with the mean field, thereby significantly reducing computational complexity. Moreover, when the number of agents is large, this approximation maintains a high degree of accuracy. Given the intricate operational and strategical details involving the platform, passengers, and drivers, constructing an exact model is exceedingly difficult, and precise analysis of such a complex model is nearly impossible. Therefore, we focus on a fundamental mean field game model that captures the core issues. Adopting an agent-based modeling perspective, we consider a closed queueing network comprising a large number of drivers, where each driver’s sojourn time follows a general distribution. The game model we propose is a continuous-time mean field game with finite state-action space where each driver’s decision model is a semi-Markov decision model (semi-MDP). A key idea for analyzing our mean field game is inspired by the analysis of closed queueing networks in (Kelly, 1989), where the mean sojourn time of a job at a queue is expressed as a function of the queue’s throughput. In the context of ride-hailing, this throughput corresponds to the driver supply rate, defined as the number of available ride-hailing vehicles per unit time within a specified region. Leveraging this insight, we transform the problem of finding mean field game equilibria into identifying the Karush-Kuhn-Tucker (KKT) points of a constrained optimization problem. This pivotal approach, originally applied to mean field games in (Courcoubetis and Dimakis, 2023), is further generalized and extended in our work.
@@
Employing the aforementioned mean field game model, this study examines the impact of customer abandonment behavior on platform dispatch algorithms. Specifically, we delve into the classical nearest-neighbor dispatch algorithm, which pairs the closest available vehicle with a customer in need. Our analysis reveals the potential for WGC. This manifests as the existence of multiple equilibria within the system, one of them being an equilibrium with longer pick-up times compared to the other equilibria. In this equilibrium, distant vehicles are matched with passengers, leading to substantial efficiency losses in the system, which theoretically could become infinitely large. Meanwhile, our analysis reveals that multiple equilibria emerge because the mean pick-up time is a non-monotone function of the driver supply rate in the given region.
@@
To prevent the emergence of such long pick-up time equilibria, we investigate the effects of controlling the matching radius – the maximum distance within which a match is allowed. Our findings indicate that the careful selection of the matching radius is crucial, as an improper choice may result in multiple equilibria exhibiting inefficiency. In light of the difficulties associated with analytically determining the optimal matching radius, our study proposes the Two-Matching-Radius Nearest-Neighbor Dispatch Algorithm, which can be viewed as a generalization of the traditional nearest-neighbor dispatch algorithm. This novel approach designs distinct matching radii for two scenarios: the arrival of new customers and the arrival of new available vehicles. Rather than seeking static optimal values for a fixed system, we derive an explicit relationship between the optimal matching radii and the driver supply rate. Furthermore, we demonstrate that the nearest-neighbor dispatch algorithm utilizing the optimal adjustment of matching radii results in a sojourn time that is a monotone function of the driver supply rate. This allows us to prove the uniqueness of the mean field equilibrium, eliminating any other inefficient equilibrium. By minimizing the sojourn time for each given driver supply rate through our optimal matching radius, our solution, besides eradicating undesirable equilibria, also maximizes the system’s throughput, thereby optimizing overall system efficiency for the ride-hailing platform.
@@
  • We introduce a mean field game model to holistically analyze the ride-hailing systems, incorporating key aspects such as driver’s repositioning across multiple regions, customer’s abandonment behavior, and platform’s dispatch algorithms. This unified model allows for the study of their joint effects on system performance.
  • In this general multi-region mean field game framework, we identify all possible equilibria as the Karush-Kuhn-Tucker (KKT) points of an associated optimization problem. Our model reveals the emergence of multiple equilibria, some of them exhibiting large pick-up times – the ‘Wild Goose Chase’ behavior in ride-hailing systems extensively discussed in the literature (Castillo et al., 2023). In such an equilibrium, drivers are matched with distant requests and consequently the system efficiency is compromised.
  • We relate for the first time the choice of the matching radius with generating multiple system-wide equilibria. We propose a simple yet innovative ‘two-matching-radius nearest-neighbor’ dispatch algorithm. This algorithm ensures a unique mean field equilibrium for the multi-region model. This implies a determinate behaviour for the whole system by eliminating multiple equilibria, some of them inefficient.
@@ 1.1. Literature Review
Repositioning in Ride-hailing Platforms. (Braverman et al., 2019) studied the repositioning of empty cars in a spatial ride-sharing network where a platform controls the repositioning of all vehicles. (Afèche et al., 2023) studied strategical repositioning in ride-hailing networks, where drivers are allowed to make repositioning decisions to maximize their earnings. (Benjaafar et al., 2021) and (Gao et al., 2024) considered a combination of decentralized and centralized repositioning to investigate the impact of autonomous vehicles on ride-hailing platforms. Our work is closely related to (Gao et al., 2024) in that we similarly consider a mean field game model and characterize the mean field equilibrium by an optimization problem. In the above literature, the matching procedure is simplified assuming a fluid model. It explicitly or implicitly assumes that rider requests are lost if not matched instantly with an available driver. In our model, both riders and drivers wait, and waiting riders abandon the system at a given rate. Our model includes a detailed matching procedure between riders and drivers, which affects the strategic repositioning behavior of drivers.
@@
Ride-matching and Abandonment. Another stream of research close to our work is the study of the performance of matching in ride-hailing systems. (Besbes et al., 2022) considers nearest-neighbor dispatch and uses a spatial queueing model to derive the service rate as a function of pick time and en-route time. (Wang et al., 2022) extends the spatial model in (Besbes et al., 2022) to include admission control for matching based on a pickup-time threshold. In addition, abandonment and cancellation on the customer’s side are investigated in (Wang et al., 2022). Our model inherits the customer’s abandonment behavior model from (Wang et al., 2022), and considers a two-sided queueing model with an open customer-side and a closed driver-side queueing network. (Castillo et al., 2023) investigated the ‘Wild Goose Chase’ caused by long pick-up times. (Castillo et al., 2023) suggested that setting a maximum matching radius could improve the system performance. (Yang et al., 2020) studied the impact of matching radius and matching time interval on the system performance. (Wang et al., 2022) and (Yang et al., 2020) both considered batch matching while (Wang et al., 2022) proposed to manipulate the matching radius and matching time interval indirectly through a pickup-time threshold. Unlike (Wang et al., 2022) and (Yang et al., 2020), we focus on instant matching using a nearest-neighbor dispatch approach. Specifically, an arriving customer is immediately matched with the closest available idle driver if one is present, and an arriving driver is immediately matched with the closest waiting customer if one is available. Thus, our work views matching radius as the key decision variable of the platform. Our research contributes to this stream of research by introducing a novel two-matching-radius nearest-neighbor dispatch algorithm. The dynamical optimal adjustment of the matching radii proposed in this work offers a solution that mitigates the WGC problem.
@@
Mean Field Game. Another area closely related to our work is mean-field games, which investigate the repeated interactions among a large number of strategic agents ((Jovanovic and Rosenthal, 1988; Huang et al., 2006; Lasry and Lions, 2007; Hopenhayn, 1992; Adlakha et al., 2015)). Our mean-field game model is of finite type, featuring finite state and action spaces. Existing finite mean-field game models in the literature are typically based on either discrete-time Markov Decision Processes (MDPs) (Guo et al., 2023) or continuous-time MDPs with exponentially distributed sojourn times (Neumann, 2020). In our model, we model each driver’s decision-making by continuous-time semi-Markov decision processes (SMDPs) where sojourn times follow general distributions. In fact, our model incorporates a closed queueing network, where drivers can be viewed as jobs navigating through the network. To analyze the closed queueing network, we adopt the approach proposed in (Kelly, 1997), which represents mean sojourn times as functions of throughputs (corresponding to driver supply rates in the ride-hailing model). This approach enables us to determine the throughputs within the closed queueing network. Building on this foundation, we utilize the methodology outlined in (Courcoubetis and Dimakis, 2023) to formulate an optimization problem—specifically, the minimization of a potential function—to compute mean-field equilibria. However, the complexity of the closed spatial queueing network inherent in ride-hailing systems often results in sojourn times that are non-monotone functions of throughput, thereby violating the monotonicity assumptions presented in (Kelly, 1989) and (Courcoubetis and Dimakis, 2023). To address this challenge, we extend their methodologies to accommodate non-monotone cases. Furthermore, by optimally designing the dispatch algorithm and matching radius, we can ensure that sojourn times exhibit monotonicity, thereby satisfying the assumptions required in (Kelly, 1989) and (Courcoubetis and Dimakis, 2023).
@@ 1.2. Paper Organization
In Section 2, we describe each driver’s continuous-time semi-Markov Decision Process model and formulate a mean field game model. In Section 3, we investigate the ride-matching algorithm under abandonment behavior, and derive the sojourn times as exact functions of driver supply rates. In Section 4, we provide the optimal matching radius calculation that eliminates multiple equilibria. In Section 5, we conduct simulations to evaluate the performance of our proposed dispatch algorithm. We conclude the paper in Section 6.
@@ 2.1. A Semi-Markov Decision Process
We consider a closed network consisting of a total mass $m$ of drivers operating across $N$ regions, labeled $1,2,\dots,N$. These regions represent distinct geographic areas within a city. Driver behavior is modeled as a continuous-time semi-Markov decision process (semi-MDP) with state space $S={1,2,\dots,N}$ the set of regions. The state of the MDP represents the region in which a driver just deposited a customer or has arrived without one. The action space, denoted by $A={1,2,\dots,N}$, represents the driver’s choice of seeking their next customer. Since both states and actions are indexed by regions, we have $S=A$.
For a given state-action pair $(i,k)\in S\times A$, the choice $k=i$ indicates that the driver remains in region $i$ until a next customer is found locally, whereas $k\neq i$ corresponds to a ‘repositioning’ decision: the driver travels empty to region $k$, where they will make a new decision upon arrival.
@@
Transition Probability. The transition probabilities are encapsulated in the tensor $P=(P^{k}_{ij})_{k\in A\atop i,j\in S}$, where $P^{k}_{ij}$ denotes the probability of transitioning from state $i$ to state $j$ under action $k$. Let $(q_{ij})_{i,j\in S}$ denote the ‘demand matrix’, where $q_{ij}$ is the probability that a passenger in region $i$ has a destination in region $j$. Using the demand matrix, we are able to define the transition matrix of the semi-MDP as follows.
  • $P_{ij}^{k},k\not=i$. The action of the driver is to reposition (without customer) to region $k\not=i$. Hence, $P_{ij}^{k}=\delta_{jk}$, where $\delta_{jk}=1$ if $j=k$ and 00 otherwise.
  • $P_{ij}^{k},k=i$. The action of the driver is to seek a customer in the same region he currently is, and this customer will dictate his next region where he will be free. Hence, $P_{ij}^{i}=q_{ij}$.
Customer Waiting. For each region $i$, customers arrive at a rate of $b_{i}>0$ and wait in queues if they are not matched with drivers upon arrival. Waiting customers abandon the system at a rate $\theta_{i}$ specific to each region. A detailed analysis of customers’ abandonment behavior is provided in Section 3 and is not required to derive the results in this section; instead, we only rely on the general structural properties of the drivers’ sojourn times to characterize the equilibria of the game.
@@
Sojourn Times. For a state-action pair $(i,k)$, it takes a certain amount of time for a driver to complete action $k$ when they are in state/location $i$ (to serve a customer in region $i$ when $k=i$ or reposition to $k$ if $k\not=i$ ). This is also referred to as the sojourn time for the state-action pair $(i,k)$. Let $T_{ik}$ denote the mean time taken for a driver to finish action $k$ from state/location $i$. The sojourn time is a result of interactions across the platform, customers, and drivers, encompassing elements such as the driver’s wait time until a match, the duration of pick-ups, and travel times between locations. Factors influencing the determination of sojourn time include customer behavior, platform matching policies, and drivers’ strategies. Under appropriate assumptions and within specific customer behavior models and platform matching policies, we will show that the mean sojourn time can be calculated based on the distribution of driver masses over the state-action space and their respective mass rates.
@@
Steady State Analysis. We consider a non-atomic game model where a representative driver makes decisions against the mass behavior of all other drivers, characterized by the mass distribution over $S\times A$. This representative driver is negligible and exerts no influence on the system, including the mass distribution and corresponding mass rate. This approach is reasonable given the enormous number of drivers in the system, their anonymity, symmetry, and the sole differentiation based on their locations. We focus on stationary equilibria and study stationary systems, which is sensible since each driver perceives the system to be in a steady state, enabling observation of mass distributions and estimation of sojourn times effectively. Our approach involves a mean field based analysis that operates with the averages of the sojourn times and associated metrics, a methodology commonly employed in related literature (Braverman et al., 2019; Besbes et al., 2022; Wang et al., 2022; Courcoubetis and Dimakis, 2023).
@@
We now determine the sojourn times in a stationary system, where all quantities refer to their expected (mean) values. Let $t_{ij}$ denote the expected trip duration from region $i$ to region $j$; in particular, $t_{ii}$ represents the expected duration of a trip within region $i$.
Consider any action $(i,k)\in S\times A$. If $k\neq i$, the vehicle performs an empty repositioning trip from region $i$ to region $k$, and the sojourn time is simply the travel time: $T_{ik}=t_{ik}$.
If $k=i$, the vehicle remains in region $i$ to serve a customer, and the sojourn time consists of three components:
  1. (1)
    the expected waiting time until the driver is matched to a customer, denoted $w_{i}^{d}$;
  2. (2)
    the expected pick-up time, denoted $\tau_{i}$; and
  3. (3)
    the expected trip time, denoted $t_{i}$.
The trip time $t_{i}$ is computed as the average over all possible customer destinations, weighted by the demand probabilities:
$t_{i}=\sum_{j\in S}t_{ij}q_{ij}.$
Note that both the driver’s waiting time $w_{i}^{d}$ and the pick-up time $\tau_{i}$ are affected by the customer arrival process and the platform’s matching policy. The relationships between the driver’s average waiting time, pick-up time, and mass rate can be established in each region based on the customer behavior model and the platform’s matching policy. We will investigate these relationships in detail in subsequent sections. However, in this section, rather than deriving their exact forms, we focus on their structural properties, which are sufficient for characterizing and proving general equilibrium properties. These properties should hold in general ridehailing scenarios when customers are impatient, and we will formally establish them within our modeling framework in subsequent sections.
@@
For a given vector of customer arrival rates and platform matching policy, when the system is in a stationary state, the following holds for every $i\in S$:
  1. (1)
    The driver’s mean waiting time $w_{i}$ and mean pick-up time $\tau_{i}$ are positive and can be written as functions of the mass rate of available drivers $x_{ii}$. Consequently, the mean total sojourn time $T_{ii}$ can also be expressed as:
@@
    where $t_{i}=\sum_{j\in S}t_{ij}q_{ij}$ is a constant.
  2. (2)
    As $x_{ii}\rightarrow b_{i}$ (the demand-supply limit), the waiting time grows unboundedly:
    $w_{i}^{d}(x_{ii})\rightarrow\infty.$
The idea of expressing $T_{ii}$ as a function of throughput ($x_{ii}$) stems from closed queuing network analysis (Kelly, 1989). In Section 3, we will demonstrate that this assumption holds for our matching mechanism and customer abandonment model, providing explicit formulae for validation. The condition $w_{i}^{d}(x_{ii})\rightarrow\infty$ as $x_{ii}\rightarrow b_{i}$ arises from customer impatience: in order to serve all customers, given that customers abandon the service at a positive rate, we need an unbounded idle driver pool.
For notational uniformity, we write the sojourn time as $T_{ik}(x)$ for all $(i,k)\in S\times A$, even though it depends only on $x_{ik}$ when $k=i$ and remains constant otherwise.
Given Assumption 1, the mass rate satisfies the following equation for a stationary system,